# EP-Diffuser: An Efficient Diffusion Model for Traffic Scene Generation and Prediction via Polynomial Representations

### v1

@@ V Conclusion
Traffic scene prediction remains a fundamental challenge in autonomous driving due to its inherently multi-modal nature. Most motion prediction competitions aim for prediction accuracy, encouraging models to focus on reproducing observed behaviors rather than capturing the diversity of plausible future scene evolutions. In this work, we introduced EP-Diffuser, a novel diffusion-based framework that leverages polynomial representations to efficiently model agent trajectories and road geometry. Through extensive experiments on Argoverse 2 and Waymo Open datasets, we demonstrated that EP-Diffuser not only improves the plausibility of generated traffic scenes but also generalizes well to out-of-distribution (OoD) environments. Notably, EP-Diffuser’s computational efficiency addresses a fundamental limitation of diffusion models, making it viable for real-time applications.
@@
Appendix A Implementation Details of Our Model
@@ A-A Diffusion and Denoising
We perform diffusion-denoising on the polynomial parameters (control points) representing each agent’s future trajectory. Specifically, we use the displacement vectors between control points. Since we employ the 6-degree Bernstein polynomials to represent 2D future trajectory data, the displacement vectors are denoted as for each agent.
Following the practice of Denoising Diffusion Probabilistic Models (DDPM) [27], we denote the diffused for the -th agent at -th diffusion step as . Here, corresponds to the fitted polynomial parameters without added noise. The forward diffusion process is expressed as:
where is the noise-scheduling parameter at diffusion step and controls the diffusion process. Equivalently, we can write as a linear combination of the initial parameters and a Gaussian noise :
We apply a total of diffusion steps to gradually transition from the data distribution to the target prior distribution .
DDPM requires steps for the backward denoising process. To accelerate this process, we adopt Denoising Diffusion Implicit Models (DDIM) [20], reducing the number of denoising steps to with a step stride . The future trajectories of agents are iteratively denoised by predicting the added noise for each agent and subtracting from at each step.
@@ A-B Encoder
We employ the encoder of EP [23] and adopt the “query-centric” design of QCNet [21], adapting it to our input data representations.
@@
Similarly, for all map elements in a scene, we encode the 2D displacement vectors between control points of map elements, modeled as 3-degree polynomials, denoted as , along with the corresponding map element categories. These features are summarized to form the map condition tokens , where is the number of map elements.
train / test | model | agent kinematic | agent interaction | map | ||||||
linear | linear | angular | angular | dist. to | collision | time to | offroad | dist. to | ||
speed | acc. | speed | acc. | agents | collision | boundary | ||||
FMAE-MA [14] (ensemble) | 0.495 | 0.328 | 0.246 | 0.210 | 0.340 | 0.802 | 0.710 | 0.758 | 0.764 | |
A2 / A2 | OptTrafDiff [7] | 0.510 | 0.367 | 0.419 | 0.537 | 0.364 | 0.848 | 0.723 | 0.863 | 0.788 |
EP-Diffuser (ours) | 0.467 | 0.343 | 0.458 | 0.759 | 0.361 | 0.845 | 0.709 | 0.863 | 0.773 | |
FMAE-MA [14] (ensemble) | 0.440 | 0.356 | 0.171 | 0.119 | 0.482 | 0.834 | 0.658 | 0.731 | 0.703 | |
A2 / WO | OptTrafDiff [7] | 0.460 | 0.374 | 0.364 | 0.416 | 0.509 | 0.905 | 0.696 | 0.878 | 0.740 |
EP-Diffuser (ours) | 0.461 | 0.391 | 0.389 | 0.584 | 0.512 | 0.921 | 0.704 | 0.900 | 0.740 |
@@ A-D Training Loss
Following DDPM [27], the EP-Diffuser is trained to minimize the mean squared error (MSE) between the added noise and predicted noise averaged across all agents:
@@ A-E Training Setup
We report the training setup for EP-Diffuser in Table IV. The noise scheduling parameter is expressed as , where .
@@ A-F Post-processing
We observe that stationary agents in recorded ground truth often exhibit minor positional shifts and unrealistic rotations, which can propagate into traffic scene prediction models, causing unnatural behaviors. This issue affects all evaluated models and is not specific to EP-Diffuser. Inspired by prior studies that apply physical constraints for motion stabilization in trajectory prediction [29, 30], we introduce a lightweight post-processing step to improve the realism of predicted traffic scenes:

### v3

@@ V Conclusion
Traffic scene prediction constitutes a fundamental challenge in autonomous driving due to its inherently multi-modal nature. Most motion prediction competitions aim for prediction accuracy, encouraging models to focus on reproducing observed behaviors rather than capturing the diversity of plausible future scene evolutions. In this work, we introduced EP-Diffuser, a novel diffusion-based framework that leverages polynomial representations to efficiently model agent trajectories and road geometry. Through extensive experiments on Argoverse 2 and Waymo Open datasets, we demonstrated that EP-Diffuser not only improves the plausibility of generated traffic scenes but also generalizes well to out-of-distribution (OoD) environments. Notably, EP-Diffuser’s computational efficiency addresses a fundamental limitation of diffusion models, making it viable for real-time applications.
@@
ACKNOWLEDGMENT
The authors would like to thank Jonas Neuhoefer and Andreas Philipp for many helpful discussions. We also gratefully acknowledge the Gauss Center for Supercomputing for providing compute on JUWELS at JSC. This work is funded by the German Federal Ministry for Economic Affairs and Energy within the project “NXT GEN AI METHODS”.
Appendix A Implementation Details
@@ A-A Diffusion and Denoising
We perform diffusion-denoising on the polynomial parameters (control points) that represent each agent’s future trajectory. Specifically, we operate on the displacement vectors between consecutive control points of 6-degree Bernstein polynomials representing 2D future trajectories, denoted as for each agent.
Following the practice of Denoising Diffusion Probabilistic Models (DDPM) [39], we denote the diffused for the -th agent at -th diffusion step as . Here, corresponds to the fitted polynomial parameters without added noise. We can write as a linear combination of the initial parameters and a Gaussian noise :
where is the noise-scheduling parameter at diffusion step and controls the diffusion process.
We apply a total of diffusion steps to gradually transition from the data distribution to the target prior distribution . For the backward denoising process, we adopt Denoising Diffusion Implicit Models (DDIM) [30] – the same procdedure as OptTrajDiff – with 10 denosing steps. The future trajectories of agents are iteratively denoised by predicting the added noise for each agent and subtracting from at each step.
@@ A-B Encoder
We employ the encoder architecture of our EP model [20] and adopt the “query-centric” design of QCNet [31], adapting it to operate on polynomial representations.
@@
Similarly, for all map elements in a scene, we encode their control point vectors of 3-degree polynomials, denoted as , along with the corresponding map element categories. These features are summarized to form the map condition tokens , where is the number of map elements.
@@ A-D Training Loss
Following DDPM [39], the EP-Diffuser is trained to minimize the mean squared error (MSE) between the added noise and predicted noise averaged across all agents: .
@@ A-E Training Setup
We report the training setup for EP-Diffuser in Table III. The noise scheduling parameter is expressed as , where .
@@ A-F Post-processing
We observe that stationary agents in recorded ground truth often exhibit minor positional shifts and unrealistic rotations, which can lead to unnatural behaviors in predicted scenes across all models. To address this, we adopt a lightweight post-processing step inspired by prior work [40]: